OpenStudy (anonymous):

How many zeros does the function f(x)=-3x^5+2x^3+7x-5 has?

OpenStudy (anonymous):

f(x)=-3x^5+2x^3+7x-5 0 =-3x^5+2x^3+7x-5 factor out an x 0 =x(-3x^4+2x^2+7)-5 now do you see how the expression in the ( ) is reducable into a quadratic? let x^2 = u

OpenStudy (anonymous):

5...tha max.power of variable gives no of zeross..oky..?

OpenStudy (anonymous):

the solution is 3, not 5, by the way

OpenStudy (anonymous):

I was thinking 4.. but am not absolutely sure

OpenStudy (anonymous):

0 =x(-3x^4+2x^2+7)-5 u = x^2 0 =x (-3u^2+2u+7)-5

OpenStudy (anonymous):

there are 3 real solutions, but also there are 2 imaginary solutions

OpenStudy (anonymous):

but the total of 5 sol to which am talking abut.,,,

OpenStudy (anonymous):

yes yes, you are correct zaibali.qasmi, i was wrong

OpenStudy (anonymous):

I only considered reals solutions as possibilities for zeros